# Semiparametric KSD test: unifying score and distance-based approaches for goodness-of-fit testing

### v1

@@ Abstract
Goodness-of-fit (GoF) tests are fundamental tools for assessing model adequacy. Score-based GoF tests are particularly appealing because they require fitting the model only once under the null. However, extending these tests to powerful nonparametric settings is challenging, due to the lack of suitable score functions. Through a class of exponentially tilted models, we show that the resulting score-based GoF tests are equivalent to the tests based on integral probability metrics (IPMs) indexed by a function class. When the class is rich, the test is universally consistent. This simple yet insightful perspective enables reinterpretation of classical distance-based testing procedures—including those based on Kolmogorov–Smirnov distance, Wasserstein-1 distance, and maximum mean discrepancy—as arising from score-based constructions. Building on this insight, we propose a new nonparametric score-based GoF test through a special class of IPM induced by kernelized Stein function class, called semiparametric kernelized Stein discrepancy (SKSD) test. Compared with other nonparametric score-based tests, the SKSD test is computationally efficient and accommodates general nuisance-parameter estimators, supported by a generic parametric bootstrap procedure. The SKSD test is universally consistent and attains Pitman efficiency. Moreover, SKSD test provides simple GoF tests for models with intractable likelihoods but tractable scores with the help of Stein’s identity and we use two popular models, kernel exponential family and conditional Gaussian models, to illustrate the power of our method. Our method achieves power comparable to task-specific normality tests such as Anderson-Darling and Lilliefors, despite being designed for general nonparametric alternatives.
@@ 1 Introduction
A classical approach to goodness-of-fit (GoF) testing is the likelihood ratio test (LRT) [wilks1938large, lehmann2005testing]. However, the LRT can be suboptimal—or even powerless—when alternatives lie outside the assumed parametric family [cox1979theoretical], particularly in the presence of more nuanced nonparametric alternatives. To address this limitation, various extensions have been proposed, most notably the generalized likelihood ratio test [perlman1999emperor, fan2001generalized, fan2005nonparametric, fan2007nonparametric]. Another line of work, which does not directly rely on the likelihood function, constructs nonparametric GoF tests by comparing the estimated distributions under the null and alternative hypotheses using distance measures such as $L_{1}$, $L_{2}$, or $L_{\infty}$ [bickel1973some, azzalini1989use, hardle1993comparing]. Both approaches typically require smoothing techniques or local estimators to capture general alternatives, but as a result, they suffer from the curse of dimensionality [gonzalez2013updated].
In light of these challenges, the celebrated score-based test [rao1948large, rao1973linear] offers an appealing alternative. A key advantage of score-based tests is that they avoid fitting the unrestricted (full) model, thereby sidestepping the need for estimation under a nonparametric specification. Surprisingly, despite their wide use in low-dimensional parametric settings, relatively little attention has been paid to their extension to nonparametric frameworks. A central obstacle is that, in fully nonparametric settings, even defining an appropriate score is nontrivial, making both its construction and the interpretation of the resulting test statistic particularly challenging.
@@ 1.1 Our contributions
Under a class of general exponentially tilted models (ETMs), we show that the score under the model is equivalent to a special class of distance measures, known as integral probability metrics (IPMs). IPMs are characterized by a class of test functions $\mathcal{F}$. This powerful connection reveals that GoF tests based on IPMs are actually score-based tests under the ETM when the model is indexed by the same function class $\mathcal{F}$. Therefore, depending on the chosen $\mathcal{F}$, several celebrated nonparametric distance-based tests—such as those based on the Kolmogorov-Smirnov (K-S) [durbin1975kolmogorov, margolin1976tests], Wasserstein-1 (W-1) distances [hallin2021multivariate] and maximum mean discrepancy [key2025composite]—can be reinterpreted from a nonparametric score-based perspective. Despite its simplicity, this connection has not been explicitly recognized in the existing literature. Such duality of score-based and distance-based tests offers a unified interpretation for seemingly different GoF testing procedures and may inspire new test procedures with appealing computational and statistical properties.
A new nonparametric score-based test.
Leveraging such duality, we introduce a nonparametric score-based test grounded in a reproducing kernel Hilbert space (RKHS) induced by kernelized Stein functions [liu2016kernelized]. We refer to this test as the semiparametric test based on kernelized Stein discrepancy, SKSD test and its semiparametric nature arises from the nonparametric kernelized Stein function class within the tilted parametric model. In particular, SKSD test offers several desirable properties:
@@
    Computational efficiency: Because of the Stein’s identity, the use of kernelized Stein functions endows the SKSD test with a closed-form expression for its test statistic, which can be computed directly from the data without requiring numerical integration. The SKSD statistic can be computed in at most $O(n^{2})$ time, which is dimension-agnostic and can even be reduced to $O(n)$ if certain choice of kernel function is considered, making it far more scalable in practice. In contrast, test statistics based on other IPMs such as the K-S or W-1 distances typically require at least $O(n^{3})$ computation for multivariate data and are thus computationally expensive.
@@
    Flexibility of nuisance estimation: The semiparametric nature of the SKSD test accommodates general estimator $\hat{\theta}_{n}$ for the “best-fit” parameter $\theta_{0}$ under null. The test remains asymptotically valid and powerful as long as the estimator $\hat{\theta}_{n}$ satisfies a mild asymptotic linear condition. To support this generality, we develop a provably valid parametric bootstrap testing framework that applies to a broad class of estimators, including two most prominent examples: M-estimators and minimum distance estimators.
@@
    Universal power and asymptotic efficiency: Under mild regularity conditions, the SKSD test is universally powerful against any fixed alternative to the null hypothesis. Moreover, we establish the rate-optimal efficiency of the proposed test by deriving its limiting power under two local contiguous alternatives (i.e., alternatives approaching $H_{0}$ at the $O(n^{-1/2})$ rate). These results establish the Pitman efficiency of the SKSD test.
@@
    Capability to test models with intractable likelihoods: The proposed test relies only on the score function under the null, making it well-suited for models with intractable likelihoods but tractable score functions. This provides a useful diagnostic tool for complex models such as exponential family graphical models [besag1975statistical, wainwright2008graphical], kernel exponential family models [canu2006kernel], and energy-based models [grathwohl2020learning], which classical GoF test methods find challenging to handle.
Our theoretical framework combines the recent advances in variational representation of kernel-based U-statistic [fernandez2024general] and stable convergence tools developed in [hausler2015stable]. Through extensive simulation results, we demonstrate that the proposed SKSD test attains power comparable to established normality tests, including the Anderson-Darling test [anderson1952asymptotic] and Lilliefors test [lilliefors1967kolmogorov]. We apply the SKSD test to the structural detection of two popular models: kernel exponential family models [canu2006kernel] and conditional Gaussian models [arnold1999conditional], where the likelihood functions are intractable in both cases. The results show that SKSD test is able to effectively detect the true model structure in these complex settings.
@@ 1.2 Related work
Broadly speaking, the SKSD test fits into an emerging literature on nonparametric testing with kernel methods. Leveraging the kernel trick, powerful kernel-based procedures have been developed in a variety of settings, including classical two-sample testing [gretton2012kernel, sejdinovic2013equivalence, schrab2023mmd, chatterjee2025boosting], point-null goodness-of-fit testing [chwialkowski2016kernel, liu2016kernelized, balasubramanian2021optimality, hagrass2026minimax], independence testing [gretton2007kernel, gretton2010consistent, deb2020measuring, albert2022adaptive, shekhar2023permutation], and conditional testing [zhang2011kernel, huang2022kernel, chatterjee2024kernel, zhang2025doubly]. In particular, the point-null GoF test with KSD is a special case of our SKSD test when $\Theta$ is only a single point.
Our SKSD test is also closely related to recent works on goodness-of-fit testing under composite null hypotheses. [key2025composite] proposed a composite goodness-of-fit test based on MMD with a dedicated estimation procedure and test statistic, while [bruck2025distribution] studied model selection and specification testing for parametric models via linearized MMD statistics with Gaussian limits. These approaches rely on MMD-based tests and numerical integration, whereas our SKSD test provides a more efficient way to handle intractable models, allowing for flexible nuisance-parameter estimators. We also note that [key2025composite] discussed a special case of a composite KSD-based test in which the nuisance parameter estimate is a particular minimum-KSD estimator, without any theoretical guarantee on the validity. This test can be viewed as a special case of the general SKSD test framework.
@@ 1.3 Organization of the paper and notations
We begin by establishing a connection between score-based tests and IPM-based tests in Section 2. Building on this bridge, we introduce the SKSD test in Section 3, together with a generic parametric bootstrap calibration. In the same section, we also derive the asymptotic properties of the SKSD test under the null and show bootstrap consistency. In Section 4, we prove that the SKSD test is universally powerful against general alternative hypothesis and analyze its local power. Section 5 illustrates the practical utility of our method through applications to goodness-of-fit testing of normality, kernel exponential family models, and conditional Gaussian models. We conclude in Section 6 with discussion on limitations and potential future work. All proofs and additional experimental results are deferred to the Appendix. Before proceeding, we summarize the notation used throughout the paper.
@@
Suppose $\mathcal{X}$ is a Polish space, i.e., a complete and separable metric space, and let $\mathscr{B}(\mathcal{X})$ denote the Borel $\sigma$-algebra generated by the open sets of $\mathcal{X}$. Consider two probability measures $\mathcal{L}_{p},\mathcal{L}_{q}\in\mathcal{P}(\mathcal{X})$, the collection of all probability measures on $(\mathcal{X},\mathscr{B}(\mathcal{X}))$, with differentiable density functions $p$ and $q$. Let $p_{\theta}$ denote the density function of $\mathcal{L}_{\theta}$ with respect to the Lebesgue measure $\lambda$. Define the score function of $\mathcal{L}_{p}$ as $s_{p}(x)=\nabla_{x}\log p(x)$. Note that this score function is different from the score $\bm{s}_{\theta}(x)$ defined in (3) since the differentiation is with respect to $x$ instead of the parameter $\gamma$. We use the bold symbol $\bm{s}_{\theta}(x)$ to denote the latter to avoid confusion.
@@
For a vector $u,v\in\mathbb{R}^{d}$ in Euclidean space, we use $\|v\|$ to denote its Euclidean norm and $\langle u,v\rangle$ to denote the inner product. For a matrix $A\in\mathbb{R}^{d_{1}\times d_{2}}$, we use $\|A\|_{2}$ to denote its operator norm, i.e., $\|A\|_{2}=\sup_{\|x\|\leq 1}\|Ax\|$. For higher order tensors, the operator norm is defined similarly. For a matrix $A\in\mathbb{R}^{d\times d}$, we use $\mathrm{Tr}(A)$ to denote its trace, i.e., $\mathrm{Tr}(A)=\sum_{i=1}^{d}A_{ii}$. Let $\mathcal{F}_{d,m}$ be the space of functions from $\mathcal{X}\subseteq\mathbb{R}^{d}$ to $\mathbb{R}^{m}$, for integers $d,m\in\mathbb{N}$ and $\mathcal{C}^{\alpha,\beta}(A,B)$ be the space of functions $f:A\times B\to\mathbb{R}$ such that $\partial_{x}^{\alpha}\partial_{y}^{\beta}f(x,y)$ is continuous for all $x\in A$ and $y\in B$.
@@
The kernel in our setting is a measurable function $K(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$. We say a kernel is symmetric if $K(x,y)=K(y,x)$ for all $x,y\in\mathcal{X}$. We say a kernel is positive definite if for any $x_{1},\ldots,x_{n}\in\mathcal{X}$ and $c_{1},\ldots,c_{n}\in\mathbb{R}$, we have $\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}K(x_{i},x_{j})\geq 0$.
@@ 2 Bridging score-based and IPM-based tests
Building on a surprisingly simple tilting model, we establish the connection of score-based tests with integral probability metrics (IPMs). We introduce such equivalence in Section 2.1, harness the insights to define nonparametric score-based tests and discuss the computational complexity of the IPM-based tests (now also score-based tests) in Section 2.2.
@@ 2.1 Exponentially tilted model: a revisit on score-based tests
Classical score-based tests, such as the Rao’s score test [rao1948large, rao1973linear], are based on the score of the parametric model. For a parametric model $\{\mathcal{L}_{(\gamma,\theta)}:(\gamma,\theta)\in\Gamma\times\Theta\subseteq\mathbb{R}\times\mathbb{R}^{k}\}$, the score is defined as:
@@
Example 1 (Testing mean of Gaussian location shift model, [student1908probable, rao1948large]).
Let $\mathcal{L}_{(\gamma,\theta)}$ be a Gaussian model with mean $\gamma$ and variance $\theta^{2}$. When testing $\gamma=0$, the score can be computed as $\bm{s}_{\hat{\theta}_{n}}(X)=-\frac{1}{\hat{\theta}_{n}^{2}}\frac{1}{n}\sum_{i=1}^{n}X_{i}$. Then if we choose the linear function class, $\mathcal{F}\equiv\{ax+b:a\in[0,1],b\in\mathbb{R}\}$, the following holds:
@@
In other words, the score-based test for testing the mean under Gaussian model can be viewed as a test based on the departure of the first moment of data from the null model.
Example 2 (Testing exponentiality under Gamma family, [moran1951random, haywood2008distribution]).
Let $\mathcal{L}_{(\gamma,\theta)}$ be the Gamma distribution family with shape parameter $\gamma+1$ and scale parameter $1/\theta$. The model is well-defined when when $\gamma>-1$ and $\theta>0$. When $\gamma=0$, the model boils down to the exponential distribution $\{\mathcal{L}_{\theta}\in\mathcal{P}(\mathcal{X}):p_{\theta}=\exp(-\theta x),\theta>0\}$. When testing exponentiality ($\gamma=0$) under the Gamma distribution family, the score can be computed as $\bm{s}_{\hat{\theta}_{n}}(X)=\frac{1}{n}\sum_{i=1}^{n}\log X_{i}-\mathbb{E}_{\mathcal{L}_{(0,\hat{\theta}_{n})}}[\log X]$. Then if we choose the logarithmic function class, $\mathcal{F}\equiv\{a\log(x):a\in[0,1]\}$, the following holds:
@@
In other words, the score-based test can be viewed as a test based on the departure of the logarithmic moment of the data from the null model when it comes to exponential model nested in Gamma family.
It is clear that these score-based tests compare the empirical moments of the transformed data with the theoretical moments under the null model. The transformation function class $\mathcal{F}$ is highly problem-specific. Moreover, the null model $p_{\theta}$ is exponentially tilted towards the alternative indexed by the transformation function: $\mathcal{L}_{(\gamma,\theta)}\propto p_{\theta}(x)\cdot\exp(\gamma f(x))$ for $f\in\mathcal{F}$. This can be easily checked for the above two examples. Such observation motivates us to generalize the score-based tests to a broader class of models and function classes under so-called, exponentially tilted model (ETM). Given a class of measurable functions $\mathcal{F}$ and any $f\in\mathcal{F}$, consider the ETM:
@@
In Example 1, the null model $p_{\theta}$ is Gaussian with mean 0 and variance $\theta$. In Example 2, $p_{\theta}$ is an exponential distribution with scale $1/\theta$, assuming $\theta>0$. Let $\hat{\theta}_{n}$ be a consistent estimate of nuisance parameter $\theta$ under the null model. The score with respect to the parameter of interest $\gamma$ is then:
@@
The maximum of difference-in-mean in (4) is known as the integral probability metric (IPM, [muller1997integral]). Intuitively speaking, the test based on IPMs compare any moment difference of the data with the null model over all the functions in a prespecified function class $\mathcal{F}$, and different choices of $\mathcal{F}$ yield different IPMs. In this way, the well-studied score-based testing procedures can be interpreted as tests based on IPMs under specific choices of $\mathcal{F}$. It is thus natural to ask whether the reverse is also true: can we interpret any IPM-based tests as score-based tests with different choices of $\mathcal{F}$? We address this question in the next section.

### v2

@@ Abstract
Goodness-of-fit (GoF) tests are fundamental tools for assessing model adequacy. Score-based GoF tests are particularly appealing because they require fitting the model only once under the null. However, extending these tests to powerful nonparametric settings is challenging, mainly due to the lack of suitable scores. Through a class of exponentially tilted models, we show that the resulting score-based GoF tests are equivalent to the tests based on integral probability metrics (IPMs) indexed by a function class. When the class is rich, the test is universally consistent. This simple yet insightful perspective enables reinterpretation of classical distance-based testing procedures—including those based on Kolmogorov–Smirnov distance, Wasserstein-1 distance, and maximum mean discrepancy—as arising from score-based constructions. Building on this insight, we propose a new nonparametric score-based GoF test through a special class of IPM induced by kernelized Stein function class, called semiparametric kernelized Stein discrepancy (SKSD) test. Compared with other nonparametric score-based tests, the SKSD test is computationally efficient and accommodates general nuisance-parameter estimators, supported by a generic parametric bootstrap procedure. The SKSD test is universally consistent and attains Pitman efficiency. Moreover, SKSD test provides simple GoF tests for models with intractable likelihoods but tractable score functions with the help of Stein’s identity. We apply the SKSD testing framework to two widely used models of this type to demonstrate the power of our method. Our method also achieves power comparable to task-specific normality tests, such as the Anderson-Darling and Lilliefors tests, despite being designed for general goodness-of-fit problems.
@@ 1 Introduction
A classical approach to goodness-of-fit (GoF) testing is the likelihood ratio test (LRT) [97, 60]. However, the LRT can be suboptimal—or even powerless—when law $\mathcal{L}$ lies outside the assumed parametric family [24]. To address this limitation, various extensions have been proposed, most notably the generalized likelihood ratio test [75, 31, 29, 30]. Another line of work, which does not directly rely on the likelihood function, constructs nonparametric GoF tests by comparing the estimated distributions under the null and alternative hypotheses using distance measures such as $L_{1}$, $L_{2}$, or $L_{\infty}$ [14, 8, 48]. Both approaches typically require smoothing techniques or local estimators to capture general alternatives, but as a result, they suffer from the curse of dimensionality [39].
In light of these challenges, the celebrated score-based test [77, 78] offers an appealing alternative. A key advantage of score-based tests is that they avoid fitting the unrestricted (full) model, thereby sidestepping the need for estimation under a nonparametric specification. Surprisingly, despite their wide use in low-dimensional parametric settings, relatively little attention has been paid to their extension to nonparametric frameworks. A central obstacle is that, in fully nonparametric settings, even defining an appropriate score is nontrivial, which makes both the construction and interpretation of the resulting test statistic particularly challenging.
@@ 1.1 Our contributions
Under a class of general exponentially tilted models (ETMs), we show that the score under the model is equivalent to a special class of distance measures, known as integral probability metrics (IPMs, [69]). IPMs are characterized by a class of test functions $\mathcal{F}$. This powerful connection reveals that GoF tests based on IPMs are actually score-based tests under the ETMs when the model is indexed by the same function class $\mathcal{F}$. Therefore, depending on the chosen $\mathcal{F}$, several celebrated nonparametric distance-based tests—such as those based on the Kolmogorov-Smirnov (K-S) [26, 65], Wasserstein-1 (W-1) distances [47] and maximum mean discrepancy [42, 58]—can be reinterpreted from a nonparametric score-based perspective. Despite its simplicity, this connection has not been explicitly recognized in the existing literature. Such duality of score-based and distance-based tests offers a unified interpretation for seemingly different GoF testing procedures and may inspire new test procedures with appealing computational and statistical properties.
A new nonparametric score-based testing framework.
Leveraging the established duality, we introduce a nonparametric score-based test grounded in a reproducing kernel Hilbert space (RKHS) induced by kernelized Stein functions [63]. We refer to this test as the semiparametric test based on kernelized Stein discrepancy, SKSD test and its semiparametric nature arises from the nonparametric kernelized Stein function class within the tilted parametric model. In particular, SKSD test offers several desirable properties:
@@
    Computational efficiency: Because of the Stein’s identity, the use of kernelized Stein functions endows the SKSD test with a closed-form test statistic, which can be computed directly from the data without any numerical integration. The SKSD statistic can be computed in at most $O(n^{2})$ time and it can even be reduced to $O(n)$ if certain choice of kernel function is considered, making it far more scalable in practice. In contrast, test statistics based on other IPMs such as the K-S or W-1 distances typically require at least $O(n^{3})$ computation for multivariate data and are thus computationally expensive.
@@
    Flexibility of nuisance estimation: The semiparametric nature of the SKSD test accommodates general estimator $\hat{\theta}_{n}$ for the “best-fit” parameter $\theta_{0}$ under null. The test remains asymptotically valid and powerful as long as the estimator $\hat{\theta}_{n}$ satisfies a mild asymptotic linear condition. To support this level of generality, we develop a provably valid parametric bootstrap testing procedure that applies to any estimator satisfying the stated condition, including two of the most prominent examples: M-estimators and minimum distance estimators.
@@
    Universal power and asymptotic efficiency: Under mild regularity conditions, the SKSD test is universally powerful against any fixed alternative to the null hypothesis. Moreover, we establish the rate-optimal efficiency of the proposed test by deriving its limiting power under two local contiguous alternatives (i.e., alternatives approaching $H_{0}$ at the $n^{-1/2}$ rate). These results establish the Pitman efficiency [76] of the SKSD test.
@@
    Capability to test models with intractable likelihoods: The proposed test relies only on the score function under the null, making it well-suited for models with intractable likelihoods but tractable score functions. This provides a useful diagnostic tool for complex models such as exponential family graphical models [13, 95], kernel exponential family models [19], and energy-based models [41], which classical GoF testing methods find challenging to handle.
Our theoretical framework combines the recent advances in variational representation of kernel-based V-statistic [33] and stable convergence tools developed in [49]. Through extensive simulation results, we demonstrate that the proposed SKSD test attains power comparable to established normality tests, including the Anderson-Darling test [5] and Lilliefors test [61]. We apply the SKSD test to the structural detection of two popular models with intractable likelihood functions: kernel exponential family models [19] and conditional Gaussian models [6]. Our empirical results show that SKSD test is able to effectively detect the true model structure in these complex settings.
@@ 1.2 Related work
Broadly speaking, the SKSD test fits into an emerging literature on nonparametric testing with kernel methods. Leveraging the kernel trick, powerful kernel-based procedures have been developed in a variety of settings, including classical two-sample testing [42, 83, 82, 20], point-null goodness-of-fit testing [23, 63, 9, 46], independence testing [44, 45, 25, 1, 84], and conditional testing [102, 52, 21, 103]. In particular, the point-null GoF test with KSD is a special case of our SKSD test when $\Theta$ is only a single point.
Our SKSD test is also closely related to recent works on goodness-of-fit testing under composite null hypotheses. [58] proposed a provably valid composite GoF test based on maximum mean discrepancy (MMD) with a dedicated estimation procedure and test statistic, while [17] studied model selection and specification testing for parametric models based on MMD statistics. These MMD-based tests rely on numerical integration to compute the test statistic, whereas our SKSD test does not require any numerical integration, allows flexible nuisance-parameter estimators and provides a much more efficient way to test models with intractable likelihoods. We also note that [58] discussed a special case of a composite KSD-based test in which the nuisance parameter estimate is a particular minimum-KSD estimator, without any theoretical guarantee on the validity. In fact, their test can be viewed as a special case of the general SKSD test framework. We defer further discussion of the concurrent work [18] to Appendix A.6.
@@ 1.3 Organization of the paper and notations
We begin by establishing the connection between score-based tests and IPM-based tests in Section 2. Building on this bridge, we introduce the SKSD test in Section 3, together with a generic parametric bootstrap calibration procedure. In the same section, we also derive the asymptotic properties of the SKSD test under the null and establish bootstrap consistency, regardless of whether the data are generated under the null or the alternative. In Section 4, we prove that the SKSD test is universally powerful against general alternative hypothesis and analyze its local power. Section 5 illustrates the practical utility of our method through applications to GoF testing of normality, kernel exponential family models, and conditional Gaussian models. We conclude in Section 6 with discussion on limitations and potential future work. All proofs and additional experimental results are deferred to the Appendix. Before proceeding, we summarize the notation used throughout the paper.
@@
Suppose $\mathcal{X}$ is a Polish space, i.e., a complete and separable metric space, and let $\mathscr{B}(\mathcal{X})$ denote the Borel $\sigma$-algebra generated by the open sets of $\mathcal{X}$. Consider two probability measures $\mathcal{L}_{p},\mathcal{L}_{q}\in\mathcal{P}(\mathcal{X})$, the collection of all probability measures on $(\mathcal{X},\mathscr{B}(\mathcal{X}))$, with differentiable density functions $p$ and $q$. Let $p_{\theta}$ denote the density function of $\mathcal{L}_{\theta}$ with respect to the Lebesgue measure $\lambda$. Define the score function of $\mathcal{L}_{p}$ as $s_{p}(x)=\nabla_{x}\log p(x)$. Note that the score function is different from the score defined in (3) since the differentiation in the former is taken with respect to $x$ instead of the parameter. We use the bold symbol $\bm{s}_{\theta}(x)$ to denote the latter to avoid confusion.
@@
For a vector $u,v\in\mathbb{R}^{d}$ in Euclidean space, we use $\|v\|$ to denote its Euclidean norm and $\langle u,v\rangle$ to denote the inner product. For a matrix $A\in\mathbb{R}^{d_{1}\times d_{2}}$, we use $\|A\|_{2}$ to denote its operator norm, i.e., $\|A\|_{2}=\sup_{\|x\|\leq 1}\|Ax\|$. For higher order tensors, the operator norm is defined similarly. For a matrix $A\in\mathbb{R}^{d\times d}$, we use $\mathrm{Tr}(A)$ to denote its trace, i.e., $\mathrm{Tr}(A)=\sum_{j=1}^{d}A_{jj}$. Let $\mathcal{F}_{d,m}$ be the space of functions from $\mathcal{X}\subseteq\mathbb{R}^{d}$ to $\mathbb{R}^{m}$, for integers $d,m\in\mathbb{N}$ and $\mathcal{C}^{\alpha,\beta}(A,B)$ be the space of functions $f:A\times B\to\mathbb{R}$ such that $\partial_{x}^{\alpha}\partial_{y}^{\beta}f(x,y)$ is continuous for all $x\in A$ and $y\in B$.
@@
The kernel in our setting is a measurable bivariate function $K(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\to\mathbb{R}$. We say a kernel is symmetric if $K(x,y)=K(y,x)$ for all $x,y\in\mathcal{X}$. We say a kernel is positive definite if, for any $x_{1},\ldots,x_{n}\in\mathcal{X}$ and $c_{1},\ldots,c_{n}\in\mathbb{R}$, we have $\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}K(x_{i},x_{j})\geq 0$.
@@ 2 Bridging score-based and IPM-based tests
In this section, building on a surprisingly simple tilting model, we establish the connection of score-based tests with the tests based on integral probability metrics (IPMs). We introduce such equivalence in Section 2.1, harness the insights to define nonparametric score-based tests and discuss several well-established IPM-based tests (now also score-based tests) in Section 2.2.
@@ 2.1 Exponentially tilted model: a revisit on score-based tests
Classical score-based tests, such as the Rao’s score test [77, 78], are based on the score of the parametric model. For a parametric model $\{\mathcal{L}_{(\gamma,\theta)}:(\gamma,\theta)\in\Gamma\times\Theta\subseteq\mathbb{R}\times\mathbb{R}^{k}\}$, the score is defined as:
@@
Example 1 (Testing mean of Gaussian location shift model, [89, 77]).
Let $\mathcal{L}_{(\gamma,\theta)}$ be a Gaussian model with mean $\gamma$ and variance $\theta^{2}$. When testing $\gamma=0$, the score can be computed as $\bm{s}_{\hat{\theta}_{n}}(X)=-\frac{1}{\hat{\theta}_{n}^{2}}\frac{1}{n}\sum_{i=1}^{n}X_{i}$. Then if we consider the linear function class, $\mathcal{F}\equiv\{ax+b:a\in[0,1],b\in\mathbb{R}\}$, the following holds:
@@
In other words, the score-based test for testing the mean under Gaussian location model with an unknown scale parameter can be viewed as a test based on the departure of the first moment of data from the null model.
Example 2 (Testing exponentiality under Gamma family, [67, 50]).
Let $\mathcal{L}_{(\gamma,\theta)}$ be the Gamma distribution family with shape parameter $\gamma+1$ and scale parameter $1/\theta$. The model is well-defined when $\gamma>-1$ and $\theta>0$. When $\gamma=0$, the model boils down to the exponential distribution family $\{\mathcal{L}_{\theta}\in\mathcal{P}(\mathcal{X}):p_{\theta}=\exp(-\theta x),\theta>0\}$. When testing exponentiality ($\gamma=0$) under the Gamma distribution family, the score can be computed as $\bm{s}_{\hat{\theta}_{n}}(X)=\frac{1}{n}\sum_{i=1}^{n}\log X_{i}-\mathbb{E}_{\mathcal{L}_{(0,\hat{\theta}_{n})}}[\log X]$. Then if we choose the logarithmic function class, $\mathcal{F}\equiv\{a\log(x):a\in[0,1]\}$, the following holds:
@@
In other words, the score-based test can be viewed as a test based on the departure of the logarithmic moment of the data from the null model when it comes to testing exponential model nested in Gamma distribution family with unknown scale parameter.
It is clear that these score-based tests essentially compare the empirical moments of the transformed data with the population moments under the null model. The full model can be constructed by exponentially tilting the null model $p_{\theta}$ towards the alternative indexed by the transformation function: $\mathcal{L}_{(\gamma,\theta)}\propto p_{\theta}(x)\cdot\exp(\gamma f(x))$ for $f\in\mathcal{F}$, which can be easily checked for the above two examples. Moreover, the function class $\mathcal{F}$ is highly problem-specific and and depends on the direction of interest being investigated. This observation motivates us to generalize score-based tests within a broader model class, the exponentially tilted model (ETM). Given a class of measurable functions $\mathcal{F}$ and any $f\in\mathcal{F}$, consider the ETM:
@@
In Example 1, the null model $p_{\theta}$ is Gaussian with mean 0 and variance $\theta^{2}$. In Example 2, $p_{\theta}$ is an exponential distribution with scale $1/\theta$. Let $\hat{\theta}_{n}$ be a consistent estimate of nuisance parameter $\theta$ under the null model. The score of model (2) with respect to the parameter of interest $\gamma$ is then given by
@@
The maximum of difference-in-mean in (4) is known as the integral probability metric (IPM, [69]). Intuitively speaking, the test based on IPMs compare any moment difference of the data with the null model over all the functions in a prespecified function class $\mathcal{F}$, and different choices of $\mathcal{F}$ yield different IPMs. In this way, the well-studied score-based testing procedures in the above examples can be interpreted as tests based on IPMs under specific choices of $\mathcal{F}$. It is thus natural to ask whether the reverse is also true: can we interpret any IPM-based tests as score-based tests with different choices of $\mathcal{F}$? We address this question in the following section.